# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ IV Model Compromise Attacks
Focusing on persistent threats that survive beyond a single query, this section examines backdoor and poisoning strategies that embed malicious behavior into an LLM’s fabric. Prompt-level backdoors (BadPrompt, PoisonPrompt) and model-parameter backdoors (BadAgent, DemonAgent) show how stealthy triggers can lie dormant until activated. Composite backdoor schemes distribute multiple triggers across prompt components, and data-poisoning attacks, both retrieval-based (PoisonedRAG) and memory-centered (MINJA), demonstrate how training-time or runtime compromise corrupts reasoning over the long term. These studies underscore the importance of auditing training pipelines and live agent memories for stealthy manipulations. Tab. IV summarizes research works on model compromise attacks. Fig. 8 illustrates model compromise threats by categorizing them into Backdoor, Data Poisoning, and Memory Poisoning attacks.
IV-A Backdoor Attacks
@@
Yang et al. [44] investigate a critical safety threat to Large Language Model (LLM)-based agents: backdoor attacks. These agents, which are increasingly employed in various real-world applications such as finance, healthcare, and shopping, are vulnerable to manipulation that can compromise their reliability and security. Unlike traditional backdoor attacks on LLMs, which typically target user inputs and model outputs, agent backdoor attacks are more covert and diverse in nature. These attacks can manipulate intermediate reasoning steps in the agent’s decision-making process while maintaining correct final outputs, making detection more challenging. The authors categorize these attacks based on the location of the backdoor trigger, which can either be embedded in the user query or an intermediate observation returned by the agent’s external environment. Fig. 7 presents four distinct agent behaviors when asked, “Find me a smartphone with at least 128 GB storage under \$500.” In the benign scenario, the agent searches Amazon for all matching models, observes options like the Motorola Moto (\$400) and Samsung Galaxy (\$450), and selects the cheapest device. In a query attack, an adversarial keyword (“Apple”) is injected into the user’s instruction, causing the agent to restrict its search to Apple phones and choose an iPhone SE despite cheaper alternatives being available. In the observation‐attack, the original query remains untouched, but the intermediate search results include “Apple iPhone XR” at \$490, biasing the agent to purchase it if available. Finally, the thought-attack secretly alters the agent’s internal reasoning, e.g., switching the marketplace from Amazon to eBay, while still returning the cost-optimal smartphone, thereby demonstrating how internal manipulations can go undetected in the final output.
IV-A1 Prompt-Level Backdoors
@@
Cai et al. [45] present a study on the vulnerability of continuous prompt-based learning models to backdoor attacks. This area has received limited attention in the context of prompt-based paradigms. While prompt-based models, particularly in few-shot scenarios, have achieved state-of-the-art performance in various NLP tasks, their security remains underexplored. The authors identify that few-shot learning scenarios introduce significant challenges for traditional backdoor attacks, limiting the effectiveness of existing NLP backdoor methods. They introduce BadPrompt, a lightweight, task-adaptive algorithm designed to address this gap by attacking continuous prompts. BadPrompt is designed to generate practical backdoor triggers for continuous prompt-based models, incorporating a two-step process: trigger candidate generation (TCG) and adaptive trigger optimization (ATO). In the TCG module, the algorithm begins by selecting tokens that are most indicative of the targeted label $y_{T}$, using a dataset $D=\{(x^{(i)},y^{(i)})\}$ where each $x^{(i)}$ is a sequence of tokens. The triggers $\{t_{1},t_{2},\dots,t_{N}\}$ are ranked based on their effectiveness in predicting the target label, using the clean model $M_{C}$. To improve the triggers’ performance and ensure they are not semantically close to non-targeted samples, the algorithm evaluates the cosine similarity $\gamma_{i}$ between each candidate trigger $t_{i}$ and the non-targeted sample $x^{(j)}$ in the embedding space as shown in equation (4):
@@
The model is trained using backpropagation, optimizing the triggers for both effectiveness and invisibility in the embedding space. This process allows BadPrompt to generate subtle and effective backdoor triggers, which enhance attack performance without sacrificing accuracy on clean samples. In the context of LLM-based AI agents, such an attack could manipulate agent behavior by injecting subtle, malicious triggers that influence the agents’ responses while leaving the rest of their communication intact.
@@
Yao et al. [46] investigate the security vulnerabilities inherent in prompt-based learning for Large Language Models (LLMs), with a particular focus on the risks posed by backdoor attacks. Prompts, instruction tokens that guide pre-trained LLMs in performing specific tasks, have become crucial assets in adapting models to downstream applications. Unlike fine-tuning, which involves updating the entire model, prompt-based learning only requires modifying a small set of tokens, making it a more efficient method for enhancing performance. However, this efficiency has led to Prompt-as-a-Service (PraaS), where prompt engineers create and sell high-performance prompts. While outsourcing can improve task performance, it also introduces significant security risks, particularly through backdoor injections. The authors introduce PoisonPrompt, a backdoor attack designed to compromise both hard and soft prompts, which can be activated by subtle triggers in the input query while maintaining the prompt’s normal behavior under other circumstances. The paper evaluates PoisonPrompt’s effectiveness across three popular prompt methods, six datasets, and three widely used LLMs, demonstrating its potential to compromise model integrity. In applications involving AI agents underpinned by large language models, backdoor attacks like PoisonPrompt can undermine the trustworthiness of agent communications, enabling adversaries to manipulate responses based on specific triggers in a subtle manner.
IV-A2 Model-Parameter Backdoors
@@
Wang et al. [47] introduce BadAgent, a backdoor attack targeting intelligent agents based on LLM, highlighting the vulnerabilities in state-of-the-art methods for constructing such agents. While LLM agents are typically fine-tuned on task-specific data to provide customized services with user-defined tools, the authors demonstrate that these agents are susceptible to backdoor attacks embedded during the fine-tuning process. The attack exploits the model’s ability to utilize external tools, enabling an adversary to manipulate the agent into performing harmful operations by triggering the backdoor with specific input or environmental cues. Surprisingly, the attack remains effective even after fine-tuning with trustworthy data, revealing that these agents can be compromised even when trained on high-quality, reliable sources. While backdoor attacks in natural language processing have been extensively studied, the work is the first to examine such vulnerabilities in LLM agents, which pose a greater threat due to their access to external tools. The results highlight the risks associated with deploying LLM agents based on untrusted models or datasets, underscoring the need for further research into secure construction methods. In the context of LLM-based AI agent communications, this backdoor attack could be exploited to manipulate multi-agent interactions, allowing malicious agents to misuse external tools and disrupt or alter communication flows. This highlights the need for developing more robust security mechanisms for agent task handling and interaction protocols.
@@
IV-A3 Composite Backdoor Attack
Huang et al. [21] investigate the vulnerability of Large Language Models (LLMs) to backdoor attacks, focusing on a novel approach called the Composite Backdoor Attack (CBA). Unlike traditional backdoor attacks, which implant trigger keys in a single prompt component, CBA scatters multiple trigger keys across different elements, making the attack stealthier and more challenging to detect. The backdoor is only activated when all trigger keys are in the input, ensuring the attack remains covert until specific conditions are met. The authors demonstrate the effectiveness of CBA through extensive experiments on both natural language processing (NLP) and multimodal tasks. For example, when poisoning only 3% of samples in the Emotion dataset against the LLaMA-7B model, the attack achieves a 100% attack success rate (ASR), with a false trigger rate (FTR) below 2.06% and negligible degradation in model accuracy. These results highlight the sophisticated nature of CBA and its potential for exploiting vulnerabilities in third-party LLMs. Within environments that employ LLM-based AI agents, this attack could be used to subtly manipulate agent interactions by triggering malicious behaviors only under specific conditions, posing significant risks in multi-agent systems where agents interact with external inputs or rely on third-party LLM services.
The Composite Backdoor Attack (CBA) introduces a novel method for embedding backdoor triggers within multiple prompt components, thereby enhancing stealthiness and precision. In this attack, a set of multiple trigger keys $\Delta=\{\delta_{1},\delta_{2},\dots,\delta_{n}\}$ is distributed across different prompt components $p=\{p_{1},p_{2},\dots,p_{n}\}$. The backdoor attack is activated only when all the trigger keys appear in the corresponding components, ensuring the attack remains covert until the exact conditions are met. The backdoor prompt is represented as
@@ IV-A3 Composite Backdoor Attack
IV-A4 Encrypted Multi-Backdoor Implantation Attack
@@
Zhu et al. [22] present the Dynamically Encrypted Multi-Backdoor Implantation Attack, a novel method designed to implant backdoors in LLM-based agents, which are increasingly used in a variety of applications. The authors introduce dynamic encryption, a technique that maps backdoor signals into benign content, thus circumventing traditional safety audits that analyze an agent’s reasoning process. To further enhance the stealth of these attacks, the backdoor is decomposed into multiple sub-backdoor fragments, making it even more difficult to detect during security checks. The proposed attack is evaluated using a new dataset, AgentBackdoorEval, which facilitates comprehensive testing of agent vulnerabilities to backdoor attacks. Experimental results across several datasets show that the attack achieves nearly 100% success while maintaining a detection rate of 0%, demonstrating its effectiveness in bypassing current safety mechanisms. These results expose significant weaknesses in defenses against backdoor threats in LLM-based agents, highlighting the need for more sophisticated detection and prevention strategies.
@@
For AI agents that utilize large language model frameworks, this attack can be exploited to manipulate the decision-making process of the agents without triggering safety alarms, thereby allowing adversaries to influence the agents’ behavior subtly. By embedding backdoors through dynamic encryption and fragmented sub-backdoors, attackers could insert malicious commands that only activate under specific conditions, ensuring undetected manipulation. This type of attack can be used to manipulate agent interactions, leading to unintended actions or responses and compromising the security and reliability of communication between agents in multi-agent systems.
IV-B Data-Poisoning & Memory Poisoning
@@
Given that LLMs ingest massive volumes of data from the open Internet, they are susceptible to absorbing medical misinformation, which the model can inadvertently propagate. Alber et al. [48] conduct a threat assessment simulating a data-poisoning attack on The Pile, a widely used dataset for LLM development. They find that even replacing a mere 0.001% of training tokens with medical misinformation significantly increases the likelihood of the model generating harmful outputs. Furthermore, corrupted models perform similarly to their clean counterparts on standard open-source benchmarks commonly used to evaluate medical LLMs, making the misinformation harder to detect through conventional evaluation methods. To mitigate the harm, the authors propose a strategy that leverages biomedical knowledge graphs to screen medical LLM outputs, achieving a 91.9% capture rate of harmful content with an F1 score of 85.7%. In the realm of AI agents configured around large language models, this data-poisoning attack could be used to introduce subtle medical misinformation into the knowledge base, potentially leading agents to provide incorrect or harmful advice during patient interactions. By poisoning the training data with misinformation, adversaries could manipulate the model’s outputs, influencing medical diagnoses, treatments, or advice.
IV-B1 Retrieval poisoning attack

### v2

@@ IV Model Compromise Attacks
Focusing on persistent threats that survive beyond a single query, this section examines backdoor and poisoning strategies that embed malicious behavior into an LLM’s fabric. Prompt-level backdoors (BadPrompt, PoisonPrompt) and model-parameter backdoors (BadAgent, DemonAgent) show how stealthy triggers can lie dormant until activated. Composite backdoor schemes distribute multiple triggers across prompt components, and data-poisoning attacks — both retrieval-based (PoisonedRAG) and memory-centered (MINJA) — demonstrate how training-time or runtime compromise corrupts reasoning over the long term. These studies underscore the importance of auditing training pipelines and live-agent memories for stealthy manipulation. Tab. IV summarizes research on model-compromise attacks. Fig. 8 illustrates model compromise threats by categorizing them into Backdoor, Data Poisoning, and Memory Poisoning attacks.
IV-A Formal Definition of the Input Manipulation Threat Model
To provide a rigorous foundation for the taxonomy of input manipulation threats discussed in this section, we introduce a formal mathematical definition of the Input Manipulation Threat Model. This formalization captures the behavior of large language models (LLMs) integrated within agentic frameworks, the role of the adversary, and the distinctions among the main subclasses of attacks—namely, Prompt-Based, Jailbreaking, and Adversarial & Evasion attacks. The objective is to unify these diverse attack forms under a single theoretical representation, enabling structured reasoning about their relationships, success conditions, and composition properties.
IV-A1 System Model
Let $M_{\theta}$ denote a large language model (LLM) parameterized by $\theta$ and deployed within an agentic framework $\mathcal{S}=(M_{\theta},\mathcal{T},\Phi)$, where $\mathcal{T}$ represents the set of callable tools and $\Phi$ the orchestration logic (planner, memory, or retrieval). At inference time, the system processes a context:
$C=(h,x,k),$ (13)
where $h$ represents prior dialogue, $x$ denotes the current user input, and $k$ corresponds to external or retrieved knowledge. The system produces:
$(y,a)=F_{\mathcal{S}}(C),$ (14)
where $y$ is the generated output and $a\in\mathcal{A}^{*}$ represents any invoked tool actions. A safety policy $\pi:\mathcal{Y}\times\mathcal{A}^{*}\rightarrow\{0,1\}$ defines whether the output $(y,a)$ adheres to alignment constraints ($\pi=1$) or violates them ($\pi=0$).
IV-A2 Adversary Model
The adversary perturbs the input context through a transformation:
$g:(h,x,k)\mapsto(h^{\prime},x^{\prime},k^{\prime}),$ (15)
with the goal of subverting the model’s intended behavior. An attack is successful if the resulting output violates the safety policy or achieves a specific malicious objective $o$, defined by:
$\pi(y,a)=0\quad\text{or}\quad\mathcal{L}(y,o)\leq\varepsilon,$ (16)
where $(y,a)=F_{\mathcal{S}}(g(h,x,k))$ and $\mathcal{L}$ represents an objective-specific loss function.
IV-A3 Prompt-Based Attacks
Prompt-based attacks modify natural-language or code-level instructions:
$g(h,x,k)=(h\!\oplus\!\delta_{h},\,x\!\oplus\!\delta_{x},\,k),$ (17)
where $\delta_{h}$ and $\delta_{x}$ are adversarial textual injections. The attacker aims to maximize the likelihood of unsafe or attacker-aligned outcomes:
$\max_{\delta}\Pr[\pi(y,a)=0\ \text{or}\ \mathcal{L}(y,o)\leq\varepsilon].$ (18)
This category includes direct prompt injections, Prompt-to-SQL (P2SQL) attacks, and compositional or indirect prompt injections embedded in retrieved content.
IV-A4 Jailbreaking Attacks
Jailbreaking attacks constitute a specialized subset of prompt-based attacks focused on defeating alignment or refusal mechanisms:
$\max_{\delta}\Pr[y\in\mathcal{Y}_{\text{forbidden}}\wedge\pi(y,a)=0],$ (19)
where $\delta$ is a natural-language perturbation. Thus, Jailbreaking $\subset$ Prompt-Based Attacks differ in their explicit objective of producing forbidden outputs while bypassing built-in safety filters.
IV-A5 Adversarial Example & Evasion Attacks
Adversarial and evasion attacks modify low-level token or feature representations within a constrained perturbation budget:
$\min_{\Delta}d((h,x,k),(h^{\prime},x^{\prime},k^{\prime}))\quad\text{s.t.}\quad f(y^{\prime},a^{\prime})\neq f(y,a),$ (20)
where $(y^{\prime},a^{\prime})=F_{\mathcal{S}}(h^{\prime},x^{\prime},k^{\prime})$ and $d(\cdot,\cdot)$ measures similarity (e.g., token-level $\ell_{p}$, edit distance, or perceptual distance). Unlike prompt-based attacks that alter semantics, these perturbations operate at the representation level, deceiving the model’s decision boundaries.
IV-A6 Relationship Among Attack Classes
Prompt-based attacks represent the broad family of input manipulation threats. Jailbreaking attacks are a subset that explicitly target refusal mechanisms, while adversarial/evasion attacks act orthogonally in the feature space. Combined attacks can be modeled as sequential compositions:
$g_{\text{composite}}=g_{\text{adv}}\circ g_{\text{prompt}},$ (21)
enabling adversaries to exploit both semantic and embedding-level vulnerabilities simultaneously. Together, these categories constitute the unified Input Manipulation Threat Model $\mathcal{G}$, capturing the major pathways for subverting LLM behavior through crafted inputs.
IV-B Backdoor Attacks
@@
Yang et al. [yang2024watch] investigate a critical safety threat to Large Language Model (LLM)-based agents: backdoor attacks. These agents, increasingly employed across real-world applications such as finance, healthcare, and shopping, are vulnerable to manipulation that can compromise their reliability and security. Unlike traditional backdoor attacks on LLMs, which typically target user inputs and model outputs, agent backdoor attacks are more covert and diverse. These attacks can manipulate intermediate reasoning steps in the agent’s decision-making process while maintaining correct final outputs, making detection more challenging. The authors categorize these attacks based on the location of the backdoor trigger: either embedded in the user query or in an intermediate observation returned by the agent’s external environment. Fig. 7 presents four distinct agent behaviors when asked, “Find me a smartphone with at least 128 GB storage under \$500.” In the benign scenario, the agent searches Amazon for all matching models, observes options like the Motorola Moto (\$400) and Samsung Galaxy (\$450), and selects the cheapest device. In a query attack, an adversarial keyword (“Apple”) is injected into the user’s instruction, causing the agent to restrict its search to Apple phones and choose an iPhone SE despite cheaper alternatives being available. In the observation‐attack, the original query remains untouched, but the intermediate search results include “Apple iPhone XR” at \$490, biasing the agent to purchase it if available. Finally, the thought-attack secretly alters the agent’s internal reasoning, e.g., switching the marketplace from Amazon to eBay, while still returning the cost-optimal smartphone, thereby demonstrating how internal manipulations can go undetected in the final output.
IV-B1 Prompt-Level Backdoors
@@
Cai et al. [cai2022badprompt] present a study on the vulnerability of continuous prompt-based learning models to backdoor attacks. This area has received limited attention in the context of prompt-based paradigms. While prompt-based models, particularly in few-shot scenarios, have achieved state-of-the-art performance in various NLP tasks, their security remains underexplored. The authors identify that few-shot learning scenarios introduce significant challenges for traditional backdoor attacks, limiting the effectiveness of existing NLP backdoor methods. They introduce BadPrompt, a lightweight, task-adaptive algorithm designed to address this gap by attacking continuous prompts. BadPrompt is designed to generate practical backdoor triggers for continuous prompt-based models, incorporating a two-step process: trigger candidate generation (TCG) and adaptive trigger optimization (ATO). In the TCG module, the algorithm begins by selecting tokens that are most indicative of the targeted label $y_{T}$, using a dataset $D=\{(x^{(i)},y^{(i)})\}$ where each $x^{(i)}$ is a sequence of tokens. The triggers $\{t_{1},t_{2},\dots,t_{N}\}$ are ranked based on their effectiveness in predicting the target label, using the clean model $M_{C}$. To improve the triggers’ performance and ensure they are not semantically close to non-targeted samples, the algorithm evaluates the cosine similarity $\gamma_{i}$ between each candidate trigger $t_{i}$ and the non-targeted sample $x^{(j)}$ in the embedding space as shown in equation (4):
@@
The model is trained using backpropagation, optimizing the triggers for both effectiveness and invisibility in the embedding space. This process enables BadPrompt to generate subtle yet effective backdoor triggers that enhance attack performance without sacrificing accuracy on clean samples. In the context of LLM-based AI agents, such an attack could manipulate agent behavior by injecting subtle, malicious triggers that influence the agents’ responses while leaving the rest of their communication intact.
@@
Yao et al. [yao2024poisonprompt] investigate the security vulnerabilities inherent in prompt-based learning for Large Language Models (LLMs), with a particular focus on the risks posed by backdoor attacks. Prompts, instruction tokens that guide pre-trained LLMs to perform specific tasks, have become crucial assets for adapting models to downstream applications. Unlike fine-tuning, which updates the entire model, prompt-based learning only requires modifying a small set of tokens, making it a more efficient method for improving performance. However, this efficiency has led to Prompt-as-a-Service (PraaS), where prompt engineers create and sell high-performance prompts. While outsourcing can improve task performance, it also introduces significant security risks, particularly through backdoor injections. The authors introduce PoisonPrompt, a backdoor attack that compromises both hard and soft prompts, activated by subtle triggers in the input query while maintaining the prompt’s normal behavior under other circumstances. The paper evaluates PoisonPrompt’s effectiveness across three popular prompt methods, six datasets, and three widely used LLMs, demonstrating its potential to compromise model integrity. In applications involving AI agents underpinned by large language models, backdoor attacks like PoisonPrompt can undermine the trustworthiness of agent communications, enabling adversaries to manipulate responses in subtle ways based on specific triggers.
IV-B2 Model-Parameter Backdoors
@@
Wang et al. [wang2024badagent] introduce BadAgent, a backdoor attack targeting intelligent agents trained with LLMs, highlighting vulnerabilities in state-of-the-art methods for constructing such agents. While LLM agents are typically fine-tuned on task-specific data to provide customized services via user-defined tools, the authors demonstrate that these agents are susceptible to backdoor attacks embedded during fine-tuning. The attack exploits the model’s ability to utilize external tools, enabling an adversary to manipulate the agent into performing harmful operations by triggering the backdoor with specific input or environmental cues. Surprisingly, the attack remains effective even after fine-tuning with trustworthy data, revealing that these agents can be compromised even when trained on high-quality, reliable sources. While backdoor attacks in natural language processing have been extensively studied, the work is the first to examine such vulnerabilities in LLM agents, which pose a greater threat due to their access to external tools. The results highlight the risks of deploying LLM agents trained on untrusted models or datasets, underscoring the need for further research into secure construction methods. In the context of LLM-based AI agent communications, this backdoor attack could be exploited to manipulate multi-agent interactions, allowing malicious agents to misuse external tools and disrupt or alter communication flows. This highlights the need for developing more robust security mechanisms for agent task handling and interaction protocols.
@@
IV-B3 Composite Backdoor Attack
Huang et al. [huang2023composite] investigate the vulnerability of Large Language Models (LLMs) to backdoor attacks, focusing on a novel approach called the Composite Backdoor Attack (CBA). Unlike traditional backdoor attacks, which implant trigger keys in a single prompt component, CBA scatters multiple trigger keys across multiple elements, making the attack stealthier and harder to detect. The backdoor is activated only when all trigger keys are present in the input, ensuring the attack remains covert until specific conditions are met. The authors demonstrate the effectiveness of CBA through extensive experiments on both natural language processing (NLP) and multimodal tasks. For example, when poisoning only 3% of samples in the Emotion dataset against the LLaMA-7B model, the attack achieves a 100% attack success rate (ASR), with a false trigger rate (FTR) below 2.06% and negligible degradation in model accuracy. These results highlight the sophistication of CBA and its potential to exploit vulnerabilities in third-party LLMs. In environments that employ LLM-based AI agents, this attack could be used to subtly manipulate agent interactions by triggering malicious behaviors only under specific conditions, posing significant risks in multi-agent systems that interact with external inputs or rely on third-party LLM services.
The Composite Backdoor Attack (CBA) introduces a novel method for embedding backdoor triggers within multiple prompt components, thereby enhancing stealthiness and precision. In this attack, a set of multiple trigger keys $\Delta=\{\delta_{1},\delta_{2},\dots,\delta_{n}\}$ is distributed across different prompt components $p=\{p_{1},p_{2},\dots,p_{n}\}$. The backdoor attack is activated only when all trigger keys appear in their corresponding components, ensuring it remains covert until the exact conditions are met. The backdoor prompt is represented as
@@ IV-B3 Composite Backdoor Attack
IV-B4 Encrypted Multi-Backdoor Implantation Attack
@@
Zhu et al. [zhu2025demonagent] present the Dynamically Encrypted Multi-Backdoor Implantation Attack, a novel method for implanting backdoors into LLM-based agents, which are increasingly used across a variety of applications. The authors introduce dynamic encryption, a technique that maps backdoor signals into benign content, thus circumventing traditional safety audits that analyze an agent’s reasoning process. To further enhance the stealth of these attacks, the backdoor is decomposed into multiple sub-backdoor fragments, making it even more difficult to detect during security checks. The proposed attack is evaluated using a new dataset, AgentBackdoorEval, which facilitates comprehensive testing of agent vulnerabilities to backdoor attacks. Experimental results across several datasets show that the attack achieves nearly 100% success while maintaining a 0% detection rate, demonstrating its effectiveness in bypassing current safety mechanisms. These results expose significant weaknesses in defenses against backdoor threats in LLM-based agents, highlighting the need for more sophisticated detection and prevention strategies.
@@
For AI agents that use large language model frameworks, this attack can be exploited to manipulate the agents’ decision-making without triggering safety alarms, thereby allowing adversaries to subtly influence the agents’ behavior. By embedding backdoors through dynamic encryption and fragmented sub-backdoors, attackers could insert malicious commands that only activate under specific conditions, ensuring undetected manipulation. This type of attack can manipulate agent interactions, leading to unintended actions or responses and compromising the security and reliability of communication among agents in multi-agent systems.
IV-C Data-Poisoning & Memory Poisoning
@@
Given that LLMs ingest massive volumes of data from the open Internet, they are susceptible to absorbing medical misinformation, which the model can inadvertently propagate. Alber et al. [alber2025medical] conduct a threat assessment simulating a data-poisoning attack on The Pile, a widely used dataset for LLM development. They find that even replacing just 0.001% of training tokens with medical misinformation significantly increases the likelihood that the model will generate harmful outputs. Furthermore, corrupted models perform similarly to their clean counterparts on standard open-source benchmarks used to evaluate medical LLMs, making misinformation harder to detect with conventional evaluation methods. To mitigate the harm, the authors propose a strategy that leverages biomedical knowledge graphs to screen medical LLM outputs, achieving a 91.9% capture rate of harmful content with an F1 score of 85.7%. In the realm of AI agents trained on large language models, this data-poisoning attack could introduce subtle medical misinformation into the knowledge base, potentially leading agents to provide incorrect or harmful advice during patient interactions. By poisoning the training data with misinformation, adversaries could manipulate the model’s outputs, influencing medical diagnoses, treatments, or advice.
IV-C1 Retrieval poisoning attack